### v1

@@ 4 Discussion
Naturally, the generality of these results needs to be established for a larger number of target combinations. We have tested two and three-element combinations of Cu, Sn, Zr, Ba and Ti targets, covering a large range of electronegativities and atomic masses. The assumption of source additivity tested for Cu-Sn was only found to be invalid when the sources were adjacent to each other. This is presumed to be due to an interaction between their plasmas, detectable as a change in the source voltage, which enhances the sputtering rate compared to single-source sputtering. However, since co-sputtering from closely-placed sources also leads to a narrower composition distribution on the substrate, this situation is less useful for materials exploration and avoiding it comes at little cost. Another likely source of problems will be elements with poor sticking coefficients (high vapour pressures), e.g. Zn. Such elements may have differing adhesion at the substrate compared to the QCM during co-sputtering, due to element interactions, or if substrate heating is used. Similarly, heavy elements could cause resputtering of co-sputtered elements in some circumstances. In such cases, the approach used here would need to be extended to incorporate these additional effects. Extending the approach to reactive sputtering is another future challenge, as the mass deposition rate detected at the QCMs includes a contribution from the gas-phase element, which cannot be distinguished in-situ without additional methods.
@@
[…]
Acknowledgements The authors thank Carl Hvarfner for helpful discussions and Younes Lablali for assisting in RBS measurements. This work is supported by Swedish Foundation for Strategic Research (SSF), the strategic research area STandUP for Energy, the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The work made use of the Myfab clean-room at Uppsala University, part of a VR and KAW funded national infrastructure, and the National Academic Infrastructure for Super computing in Sweden (NAISS). Operation of the accelerator, used for RBS measurements, is supported by the Swedish Research Council VR-RFI (Contracts 2019_00191 & 2023_00155).

### v2

@@ 4 Discussion
Naturally, the generality of these results needs to be established for a larger number of target combinations. We have tested two and three-element combinations of Cu, Sn, Zr, Ba, Ti and Ge targets, covering a large range of electronegativities and atomic masses. Some further examples of the linear additivity of multiple sources are shown in the Supplementary information Figure S2 (two sources) and Figure S3 (three sources). The assumption of source additivity was only found to be invalid when the sources were adjacent to each other, Ba-Sn in Figure S2. This is presumed to be due to an interaction between their plasmas, detectable as a change in the source voltage, which enhances the sputtering rate compared to single-source sputtering. However, since co-sputtering from closely-placed sources also leads to a narrower composition distribution on the substrate, this situation is less useful for materials exploration and avoiding it comes at little cost. Another likely source of problems will be elements with poor sticking coefficients (high vapour pressures), e.g. Zn. Such elements may have differing adhesion at the substrate compared to the QCM during co-sputtering, due to element interactions, or if substrate heating is used. Figure S2 presents additional comparisons between predicted composition and RBS measurements for three Cu-Ba samples. The agreement is again remarkable in two cases. In the third case - for a high sputter pressure - the match is poorer, while the n value for Ba also becomes extreme. This indicates a less physically-realistic solution of the flux model, i.e. a highly directional flux, that is possibly connected to poor sticking of Ba at higher pressures. High values of n may therefore be used as an indicator of irregular behaviour. Finally, heavy elements could cause resputtering of co-sputtered elements in some circumstances. This would be harder to notice using QCM data alone and would rely on ex-situ measurements of prepared samples. Extending the approach to reactive sputtering is another future challenge, as the mass deposition rate detected at the QCMs includes a contribution from the gas-phase element, which cannot be distinguished in-situ without additional methods.
@@
Acknowledgements
The authors thank Carl Hvarfner for helpful discussions and Corrado Comparotto and Younes Lablali for assisting in RBS measurements. This work is supported by Swedish Foundation for Strategic Research (SSF), the strategic research area STandUP for Energy, the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The work made use of the Myfab clean-room at Uppsala University, part of a VR and KAW funded national infrastructure, and the National Academic Infrastructure for Super computing in Sweden (NAISS). Operation of the accelerator, used for RBS measurements, is supported by the Swedish Research Council VR-RFI (Contracts 2019_00191 & 2023_00155).
Appendix A Experimental settings
The following experimental setting were used for the flux model
  • 1.
    Angle to magnetron $\varphi\colon 28.4\degree$
  • 2.
    Angle to substrate $\theta\colon 43.2\degree$
  • 3.
    Distance to substrate $r\colon 155$$m$
Figure 12 shows a scale diagram of the sputter chamber, where we see how the substrate, QCMs and magnetrons are placed, both from above and the side.
Figure 12: Sputter chamber showing the magnetron and sample placements.
Appendix B Additional multi-element verification
Figure 13 and 14 compares the expected deposition rate from taking the sum of the GPs from two respectively three different elements, with the actual deposition rate during co-sputtering. The positions of the targets in the sputtering chamber are also indicated in the figure lower right corner.
Figure 13: Comparisons of the expected deposition rates based on the sum of two trained GPs (one for each target) and the measured deposition rates obtained when the two targets are sputtered together. The dashed diagonal lines are where points should fall for an accurate prediction. The relevant targets in each figure, their positions and the position of the QCMs are shown in the inset diagrams. Deviations from the expected additive behaviour are seen when the two targets are neighbours, due to plasma-plasma interactions - see for the case of Sn-Ba.
Figure 14: Comparisons of the expected and actual deposition rates for 3 sources, Cu, Ge and Ti. Sputtered with randomized conditions over 100 runs, similar to the analysis in Figure 13
Figure 15 shows the comparison of RBS measurements and predicted compositions across 3 Ba-Cu samples. The sputter conditions, model parameters and prediction error are displayed in each figure.
Figure 15: Comparison of RBS measurements and predicted compositions across 3 Cu-Ba samples. Sputtering conditions and model parameters are indicated in the figures.